Architecture: A roofing company is preparing bids on two special jobs involving cone- shaped roofs. Roofing material is usually sold in 100-square-foot squares.
For each roof, find the lateral surface area to the nearest square foot. Then determine the squares of roofing materials that would be needed to cover each surface.
Roof # 1. The cone-shaped part is 23ft with a radius of 8ft.
Roof # 2 the cone-shape part is 12ft. with a radius of 9ft. find the answer for each roof.

Answer by stanbon(75887)

Then determine the squares of roofing materials that would be needed to cover each surface.
Roof # 1. The cone-shaped part is 23ft with a radius of 8ft.
Area = (1/3)Bh
I assume the height of the cone is 23 ft.
=(1/3)(pi)(8^2)(23)
=1541.47
-----------------------

Roof # 2 the cone-shape part is 12ft. with a radius of 9ft. find the answer for each roof.
Area = (1/3)Bh
Area = (1/3)(pi)9^2
Area = 84.82 sq. ft.
===========
Cheers,
Stan H.